# Extracting Sentence Embeddings from Pretrained Transformer Models

### v1

@@ Abstract
Background/introduction: Pre-trained transformer models shine in many natural language processing tasks and therefore are expected to bear the representation of the input sentence or text meaning. These sentence-level embeddings are also important in retrieval-augmented generation. But do commonly used plain averaging or prompt templates surface it enough?
Methods: Given 110M parameters BERT’s hidden representations from multiple layers and multiple tokens we tried various ways to extract optimal sentence representations. We tested various token aggregation and representation post-processing techniques. We also tested multiple ways of using a general Wikitext dataset to complement BERTs sentence representations. All methods were tested on 8 Semantic Textual Similarity (STS), 6 short text clustering, and 12 classification tasks. We also evaluated our representation-shaping techniques on other static models, including random token representations.
Results: Proposed representation extraction methods improved the performance on STS and clustering tasks for all models considered. Very high improvements for static token-based models, especially random embeddings for STS tasks almost reach the performance of BERT-derived representations.
Conclusions: Our work shows that for multiple tasks simple baselines with representation shaping techniques reach or even outperform more complex BERT-based models or are able to contribute to their performance.
@@ 1 Introduction
Early work on learnable word-level representations [1] showed that semantic meaning can be embedded in numerical vector representations. Arithmetic operations such as king - man + woman $\approx$ queen were valid. But can similar or higher abilities be also achieved for whole sentences, not just individual words? They would enable better clustering, classification, and other tasks depending on the whole meaning of the word sequence.
@@
At the core of recent advancements in artificial intelligence is the transformer architecture [2]. Compared to previous RNN-based models, it is fully parallelizable, allowing faster training and greater generalization acquisition from the data. On the 10th birthday, a child is expected to have already experienced the meaning of more than 100 million words [3], and now large language models surpass such scales by multiple orders of magnitude. In addition to improved throughput, transformer models are based on a self-attention mechanism [4] that allows the model to attend to relevant parts of the input sequence. That is similar to how humans understand and experience the meaning of words: using context. The success of the transformer architecture in Natural Language Processing (NLP) tasks, starting with the BERT model [5], was also repeated in other fields such as vision [6], speech [7], and reinforcement learning [8]. Transformer models enabled the solving of more sophisticated tasks such as sentiment analysis or question answering, as well as became the state-of-the-art in almost every NLP task. Therefore, it should possess the representation of the whole text sequence.
@@
Transformer models generally have at least 12 layers and their parameters are counted in over hundred millions. Condensing the sentence representation using these weights into a fixed 768-length vector (a common length for token vectors in base transformer models) is a challenge. It turns out that simply extracting features from a transformer model’s last layer activations yields even worse results than much simpler models [9]. There were multiple works that tried to optimize the architecture parameters (see [10] that evaluated multiple proposed modifications), and multiple investigative works probing the properties of different parts of the model (see a critical review [11] of such approaches). Only the following ways gave tangible results: 1) the plain embedding averaging of all tokens comprising the sequence; 2) engineering a prompt template to condensate sentence representation into a single token; and 3) using a specially dedicated fine-tuned model designed to produce such vectors.
@@
Fine-tuning is definitely the most efficient option. Resulting models, such as InferSent [12] or Sentence-BERT [9], showed state-of-the-art performance in sentence-level tasks at the time. However, the success of fine-tuning depends on several factors. It requires high quantity and quality target domain data, as well as computational resources, which may not always be available. Sometimes, even the existing target domain data cannot be used as it is very expensive to label. Other difficulties emerge if the data contain sensitive or private information and present a risk of it surfacing during inference. Having such constraints, one has to resort to the first two options of using raw encoded features to produce a vector for a text sequence.
@@
One of the two pre-training tasks of the famous BERT model [5] was the next sentence prediction. It was optimized through a special [CLS] token, which sought to capture the whole-sequence-level representation. Yet in later works such as [9] it was revealed that such a representation is very poor, not better than the classic ones, and opted for simple averaging of the last layer tokens instead. The authors of [13] proposed to use averaging of tokens of the first and the last layers, and the authors of [14] also included hidden token representations of the second layer. But is that really the best way to get a numerical representation of the text’s meaning?
Pre-trained models like BERT capture a lot of useful representations, yet it is not that trivial to extract them. The authors of [15] showed that further improvements can be achieved by removing the most frequent, sub-word, uppercase, and punctuation tokens before averaging. Furthermore, even larger gains can be achieved by using a prompt template This sentence : "[X]" means [MASK], where the target sentence is placed instead of [X] and the representation of the token [MASK] is used as the final representation of the whole sentence. Such a method is a good example of representation extraction without any specific fine-tuning.
Inspired by the above findings, we hypothesize that there may be more ways to surface relevant sentence-level representations without directly fine-tuning the pre-trained BERT model. More concretely, in our analysis, we try to find a function that would shape and extract representations of bert base-uncased along its layers, target tokens, and additional corpus, so that the best performance in multiple tasks would be achieved. We evaluated our approach on short text clustering, semantic textual similarity, and classification tasks.
Considering the difficulties of acquiring representations from the state-of-the-art transformer models and following the success of recent works surfacing it, our approach offers these main contributions:
  • We experimentally test how multiple combinations of various token aggregation and sentence representation post-processing techniques impact the performance of various tasks and properties of representations on multiple models.
@@
The rest of this paper is organized as follows. We provide a review of related work in the literature on composing word vectors and representation reshaping to get sentence- or text-level embeddings in Section 2. In Section 3, we outline the experimental setting and give a detailed background on our chosen approach, models, and datasets. In Section 4, we present the results. Finally, we summarize the findings of this work in Section 5.
@@ 2 Related work
Taking BERT model as an example, we are interested in how representation can be aggregated over tokens, layers, and possibly modified; and which models produce the best representations. Finally, we look at the evaluation options.
@@ 2.1 Composing word vectors
Thousands of works are being done on word-level vectors. Now, you can easily download popular publicly accessible word2vec [16] and GloVe [17] embeddings. They are lightweight and usually fit perfectly various word-level tasks. On the other hand, modeling sequences of words is considered much more challenging. Count-based approaches lose information on word order and are sparse. Learning higher-order n-gram vector space models, not just words but phrases or sentences, leads to sparsity, as frequencies vanish for target n-grams and contexts. The latter, in particular, is the main driving force for distributed representations. We will cover special methods dedicated to sentence level in Subsection 2.3. Nevertheless, the easiest solution is to reuse individual word vectors of the sequence.
@@
Unfortunately, formal methods have many practical shortcomings. A logic-based system must explicitly maintain the lexical knowledge necessary for the inference. Therefore, expensive human labor is involved in the construction and maintenance of the knowledge resources. More importantly, it must be domain-specific. One cannot simply use all the knowledge bases of the entire community as this would hinder complex inferences of theorem provers and model building. In fact, there is an area of research on how to reduce processing time, given the large amount of knowledge resources [23]. As a result, logic-based systems have been criticized for their lack of robustness and scalability: implemented systems tend to be small-scale and domain-specific [24]. Although being white box, theoretically clear and promising, in practice, formal semantics methods are often surpassed by unsupervised distributional approaches capable of utilizing huge amounts of data. As shown in [25], symbolic representations at least can provide additional features for neural approaches.
Tensor products
@@
In [26, 24], it was proposed to combine two representations, each in vector form, using tensor products. I.e., every element of the first vector is multiplied by every element of the second vector, retaining the products of all the pair combinations. This way, two rank-1 tensors result in a rank-2 tensor. Combining more vectors results in higher rank tensors. It allows the tensors to represent the relations and role-filler bindings in a distributed fashion. However, it raises problems due to the dimensionality growing exponentially in size as more constituents are composed [27]. Furthermore, as noted in [28], tensor-based models can only efficiently handle sentences of a fixed structure. Unfortunately, in most practical applications, this is not the case.
@@
Some methods try to solve the mentioned problems by sticking to the original vector dimensionality. It can be done using convolution methods [29, 30]. For example, the circular convolution, as presented in [31], achieves compression by summing along the transdiagonal elements of the tensor product. The compression is lossy, but the noisy version of the original vector can be recovered using circular correlation [27] and matched to the original one by comparing with all known component vectors. Due to the same mathematical principles as light holography, these models are also referred to as holographic.
@@
Some works related to tensor products stem from formal semantics in a way that syntax drives the compositional process. In particular, it was expected that one of the most important parts of the sentence is the pair of a noun and an adjective. Machine learning, in particular regression-based, can be used to find the composition method. The idea is that vectors for AN (Adjective-Noun) pairs can be learned in the same fashion as for regular words. Then, given the constituents and the final AN vector, a linear mapping is learned. In [32] the mapping is learned as a separate matrix for each adjective, while in [33] a generic “AN-slot” function is trained. These methods do not increase the dimensionality, yet it is not clear how they would work with more words than only the AN pairs. A later work [34] extended the idea of the adjective matrix of [32] to other types (in addition to adjectives and nouns). They emulate formal semantics by representing functions as tensors and arguments as vectors. This way, subjects and objects are rank-1 tensors, while verbs are rank-3. However, as the authors trained only nouns, verbs, subject-verb pairs, and subject-verb-object triplets, their method is still limited to phrases no longer than 3 words.
@@
Although tensor products share some exotic mathematical properties with quantum mechanics, for example, quantum entanglement [30], the framework relies heavily on formal semantics (see [35]). Therefore, it also shares the same weaknesses and is outperformed by pure-machine-learning approaches.
Averaging
@@
As neural vectors became more effective than traditional approaches [28], it turned out that it was common to derive a vector for a sequence of words simply by summing or averaging individual vectors. In 2009, the authors of [27] reported that for composing a phrase representation “averaging is the most common form of vector combination”. Despite its simplicity, the basic rule of averaging word vectors was shown to work very well [36, 37, 38]. Later work such as [39, 40, 41] even showed that simple pooling methods, such as basic vector addition or average, match or outperform much more sophisticated methods for encoding the meaning of a text sequence. The authors of [42] reported that in out-of-domain scenarios, simple architectures such as word averaging outperform complex Long Short-Term Memory (LSTM) models. In addition, they are even competitive with systems tuned for particular tasks. Later work by [43] evaluated various compositional models and found that word vector averaging performed reasonably well in most supervised benchmarks. According to the authors of [9], even for the currently trending BERT transformer model, to map the sentence to a single vector, the most popular approach is to average the word embeddings of the BERT output layer.
@@
As averaging is well-suited to acquire sentence optimization, some methods take averaging into their structure to further optimize it. One such example is the deep average network (DAN) [44]. It takes the input as an averaged text sequence and passes it to one or more feedforward layers to finally perform linear classification. Another, C-PHRASE model of [45] is trained to predict the contexts of phrases from the additive combination of their elements. Such design results in a useful property of C-PHRASE, that summing word vectors yields sequence representation. The authors of [46] found that one should use an average of all hidden states of LSMT, rather than use representation only from the last one. In a similar fashion, the averaging or summing is optimized in the works of Siamese CBOW model [47], Sent2Vec model [48] and in other works [49, 50, 51].
@@
A model derived from BERT, specially designed for sentence embeddings, SBERT, also uses averaging to pool words from two sequences to a pair of fixed-size sentence embeddings. Later, using the pair and its element-wise difference, the representation is passed to the softmax layer for classification and regression tasks. The authors of [52] tried to improve this configuration by replacing the averaging with a convolutional neural network. However, their improvement was limited to only better scores for the SALBERT model, which was still lower than the SBERT based on averaging. Therefore, calculating the mean embedding from multiple word vectors is still a solid approach to the inner workings of current sentence models.
@@
Averaging is used in various scenarios. Some authors employ it to derive static embeddings from contextualized models. In the work of [49], word vectors are learned by predicting them from the average of all contextual embeddings of words (except the target word) returned by the BERT encoder. Other authors of [53] derived static word embeddings from contextualized transformer representations by averaging word embeddings for each word in 100k different contexts. Such derived static versions are of better quality than classic word2vec and share their benefits, such as having tens of millions of times lower computational cost than using standard contextual embedding models [49]. The other scenario is to compose a new more expressive meta-embedding from different models and domains. As a means of composition, the authors of [54] proposed concatenation with some trainable methods. However, later work of [55] found that averaging word vectors and padding them with zeros to compensate for dimensionality mismatches is a surprisingly effective method.
@@
Despite its popularity, averaging has inherent deficiencies. The most prominent is the loss of word-order information. For this reason, it is sometimes called a bag-of-words representation (that is, imagine that words lay in the bag in no specific order). As shown in [44], models based on averaging are weaker in tasks that require syntax information. The mean pooled static representations of the words “dog chased human” and “human chased dog” will be the same. In addition, simple summation can cause destructive interference, affecting valuable information. This is especially relevant in long documents. In addition to neglected sentence structure and word order, the authors of [43] also note that individual word identities are lost and noninformative words are more prominently represented than essential ones.

### v2

@@ Abstract
Pre-trained transformer models shine in many natural language processing tasks and therefore are expected to bear the representation of the input sentence or text meaning. These sentence-level embeddings are also important in retrieval-augmented generation. But do commonly used plain averaging or prompt templates sufficiently capture and represent the underlying meaning? After providing a comprehensive review of existing sentence embedding extraction and refinement methods, we thoroughly test different combinations and our original extensions of the most promising ones on pretrained models. Namely, given 110 M parameters, BERT’s hidden representations from multiple layers, and many tokens, we try diverse ways to extract optimal sentence embeddings. We test various token aggregation and representation post-processing techniques. We also test multiple ways of using a general Wikitext dataset to complement BERT’s sentence embeddings. All methods are tested on eight Semantic Textual Similarity (STS), six short text clustering, and twelve classification tasks. We also evaluate our representation-shaping techniques on other static models, including random token representations. Proposed representation extraction methods improve the performance on STS and clustering tasks for all models considered. Very high improvements for static token-based models, especially random embeddings for STS tasks, almost reach the performance of BERT-derived representations. Our work shows that the representation-shaping techniques significantly improve sentence embeddings extracted from BERT-based and simple baseline models.
@@ 1 Introduction
Early work on learnable word-level representations [1] showed that semantic meaning can be embedded in numerical vector representations. Arithmetic operations such as king -- man + woman $\approx$ queen were valid. But can similar or higher abilities also be achieved for whole sentences, not just individual words? They would enable better clustering, classification, and other tasks depending on the whole meaning of the word sequence.
@@
At the core of recent advancements in artificial intelligence is the transformer architecture [2]. Compared to previous RNN-based models, it is fully parallelizable, allowing faster training and greater generalization acquisition from the data. On their 10th birthday, a child is expected to have already encountered and understood the meaning of more than 100 million words [3], and now large language models surpass such scales by multiple orders of magnitude. In addition to improved throughput, transformer models are based on a self-attention mechanism [4] that allows the model to attend to relevant parts of the input sequence. That is similar to how humans understand and experience the meaning of words: using context. The success of the transformer architecture in Natural Language Processing (NLP) tasks, starting with the BERT model [5], was also repeated in other fields such as vision [6], speech [7], and reinforcement learning [8]. Transformer models enabled the solving of more sophisticated tasks such as sentiment analysis or question answering, as well as became the state-of-the-art in almost every NLP task. Therefore, it should possess the representation of the whole text sequence.
@@
Transformer models generally have at least 12 layers and their parameters are counted in hundreds of millions or much more. Condensing the sentence representation using these weights into a fixed 768-length vector (a common length for token vectors in base transformer models) is a challenge. It turns out that simply extracting features from a transformer model’s last layer activations yields even worse results than much simpler models [9]. There were multiple works that tried to optimize the architecture parameters (see [10], which evaluated multiple proposed modifications), and multiple investigative works probing the properties of different parts of the model (see a critical review [11] of such approaches). Only the following methods gave tangible results: (1) the plain embedding averaging of all tokens comprising the sequence; (2) engineering a prompt template to condensate sentence representation into a single token; and (3) using a specially dedicated fine-tuned model designed to produce such vectors.
@@
Fine-tuning is definitely the most efficient option. Resulting models, such as InferSent [12] or Sentence-BERT [9], showed state-of-the-art performance in sentence-level tasks at the time. However, the success of fine-tuning depends on several factors. It requires high-quantity and high-quality target domain data, as well as computational resources, which may not always be available. Sometimes, even the existing target domain data cannot be used, as they are very expensive to label. Other difficulties emerge if the data contain sensitive or private information and present a risk of it surfacing during inference. Having such constraints, one has to resort to the first two options of using raw encoded features to produce a vector for a text sequence.
@@
One of the two pre-training tasks of the famous BERT model [5] was next sentence prediction. It was optimized through a special [CLS] token, which sought to capture the whole-sequence-level representation. But in later works, such as [9], it was revealed that such a representation is very poor, not better than the classic ones, and the authors opted for simple averaging of the last layer tokens instead. The authors of [13] proposed using averaging of tokens of the first and last layers, and the authors of [14] also included hidden token representations of the second layer. But is that really the best way to obtain a numerical representation of the text’s meaning?
Pre-trained models like BERT capture a lot of useful representations, yet it is not that trivial to extract them. The authors of [15] showed that further improvements can be achieved by removing the most frequent, sub-word, uppercase, and punctuation tokens before averaging. Furthermore, even larger gains can be achieved by using a prompt template, “This sentence: "[X]" means [MASK]”, where the target sentence is placed instead of [X] and the representation of the token [MASK] is used as the final representation of the whole sentence. Such a method is a good example of representation extraction without any specific fine-tuning.
Inspired by the above findings, we hypothesize that there may be more ways to distill relevant sentence-level embeddings without directly fine-tuning the pre-trained BERT model. More concretely, in our analysis, we try to find a function that would shape and extract representations of bert base-uncased along its layers, target tokens, and additional corpus, so that the best performance in multiple tasks would be achieved. We evaluate our approach on short text clustering, semantic textual similarity, and classification tasks.
Considering the difficulties of acquiring representations from the state-of-the-art transformer models and following the success of recent works demonstrating it, our approach offers these main contributions:
  • We provide an extensive and organized review of related work on producing sentence-level embeddings from transformer models.
  • We experimentally test how multiple combinations of various of the most promising token aggregation and sentence representation post-processing techniques impact the performance of three classes of different tasks and properties of representations on several models.
@@
The rest of this paper is organized as follows. We provide a review of related work in the literature on composing word vectors and representation reshaping to obtain sentence- or text-level embeddings in Section 2. In Section 3, we outline the experimental setting and give a detailed background on our chosen approach, models, and datasets. In Section 4, we present the results. Finally, we summarize the findings of this work in Section 5.
@@ 2 Related Work
Taking the BERT model as an example, we are interested in how representation can be aggregated over tokens, layers, and possibly modified, and which models produce the best representations. Finally, we look at the evaluation options.
@@ 2.1 Composing Word Vectors
Thousands of works are being carried out on word-level vectors. Now, you can easily download popular publicly accessible Word2Vec [16] and GloVe [17] embeddings. They are lightweight and usually perfectly fit various word-level tasks. On the other hand, modeling sequences of words is considered much more challenging. Count-based approaches lose information on word order and are sparse. Learning higher-order n-gram vector space models, not just words, but phrases or sentences, leads to sparsity, as frequencies vanish for target n-grams and contexts. The latter, in particular, is the main driving force for distributed representations. We will cover special methods dedicated to sentence level in Section 2.3. Nevertheless, the easiest solution is to reuse individual word vectors of the sequence.
@@ 2.1.1 Formal Semantics
Unfortunately, formal methods have many practical shortcomings. A logic-based system must explicitly maintain the lexical knowledge necessary for the inference. Therefore, expensive human labor is involved in the construction and maintenance of these knowledge resources. More importantly, it must be domain-specific. One cannot simply use all the knowledge bases of the entire community, as this would hinder complex inferences of theorem provers and model building. There is an area of research on reducing processing time, given the large amount of knowledge resources [23]. Due to this property, logic-based systems have been criticized for their lack of robustness and scalability; implemented systems tend to be small-scale and domain-specific [24]. Although being white-box, theoretically clear, and promising, in practice, formal semantics methods are often surpassed by unsupervised distributional approaches capable of utilizing huge amounts of data. As shown in [25], symbolic representations can at least provide additional features for neural approaches.
2.1.2 Tensor Products
@@
In [26, 24], it was proposed to combine two representations, each in vector form, using tensor products, i.e., every element of the first vector is multiplied by every element of the second vector, retaining the products of all the pair combinations. This way, two rank-1 tensors result in a rank-2 tensor. Combining more vectors results in higher-rank tensors. It allows the tensors to represent the relations and role-filler bindings in a distributed fashion. However, it raises problems due to the dimensionality growing exponentially in size as more constituents are composed [27]. Furthermore, as noted in [28], tensor-based models can only efficiently handle sentences of a fixed structure. Unfortunately, in most practical applications, this is not the case.
@@
Some methods try to solve the mentioned problems by sticking to the original vector dimensionality. It can be accomplished using convolution methods [29, 30]. For example, the circular convolution, as presented in [31], achieves compression by summing along the transdiagonal elements of the tensor product. The compression is lossy, but the noisy version of the original vector can be recovered using circular correlation [27] and matched to the original one by comparing with all known component vectors. Due to the same mathematical principles as light holography, these models are also referred to as holographic.
@@
Some works related to tensor products stem from formal semantics in a way that syntax drives the compositional process. In particular, it was expected that one of the most important parts of the sentence is the pair of a noun and an adjective. Machine learning, in particular regression-based, can be used to find the composition method. The idea is that vectors for AN (adjective–noun) pairs can be learned in the same fashion as for regular words. Then, given the constituents and the final AN vector, a linear mapping is learned. In [32], the mapping is learned as a separate matrix for each adjective, while in [33], a generic “AN-slot” function is trained. These methods do not increase the dimensionality, yet it is not clear how they would work with more words than only the AN pairs. A later work [34] extended the idea of the adjective matrix of [32] to other types (in addition to adjectives and nouns). They emulate formal semantics by representing functions as tensors and arguments as vectors. This way, subjects and objects are rank-1 tensors, while verbs are rank-3. However, as the authors trained only nouns, verbs, subject–verb pairs, and subject–verb–object triplets, their method is still limited to phrases no longer than three words.
@@
Although tensor products share some exotic mathematical properties with quantum mechanics, for example, quantum entanglement [30], the framework relies heavily on formal semantics (see [35]). Therefore, it also shares the same weaknesses and is outperformed by pure machine learning approaches.
2.1.3 Averaging
@@
As neural vectors became more effective than traditional approaches [28], it turned out that it was common to derive a vector for a sequence of words simply by summing or averaging individual vectors. In 2009, the authors of [27] reported that for composing a phrase representation, “averaging is the most common form of vector combination”. Despite its simplicity, the basic rule of averaging word vectors was shown to work very well [36, 37, 38]. Later work, such as [39, 40, 41], even showed that simple pooling methods, such as basic vector addition or averaging, match or outperform much more sophisticated methods for encoding the meaning of a text sequence. The authors of [42] reported that in out-of-domain scenarios, simple architectures such as word averaging outperform complex Long Short-Term Memory (LSTM) models. In addition, they are even competitive with systems tuned for particular tasks. Later work by [43] evaluated various compositional models and found that word vector averaging performed reasonably well in most supervised benchmarks. According to the authors of [9], even for the currently trending BERT transformer model, to map the sentence to a single vector, the most popular approach is to average the word embeddings of the BERT output layer.
@@
As averaging is well-suited to acquiring sentence optimization, some methods take averaging into their structure to further optimize it. One such example is the deep average network (DAN) [44]. It takes the input as an averaged text sequence and passes it to one or more feedforward layers to finally perform linear classification. Another, C-PHRASE model of [45], is trained to predict the contexts of phrases from the additive combination of their elements. Such a design results in a useful property of C-PHRASE, that summing word vectors yields sequence representation. The authors of [46] found that one should use an average of all hidden states of LSMT, rather than using representation only from the last one. In a similar fashion, the averaging or summing is optimized in the works of the Siamese CBOW model [47], Sent2Vec model [48], and in other works [49, 50, 51].
@@
A model derived from BERT, specially designed for sentence embeddings, SBERT, also uses averaging to pool words from two sequences to a pair of fixed-size sentence embeddings. Later, using the pair and its element-wise difference, the representation is passed to the softmax layer for classification and regression tasks. The authors of [52] tried to improve this configuration by replacing the averaging with a convolutional neural network. However, their improvement was limited to only better scores for the SALBERT model, which was still lower than SBERT based on averaging. Therefore, calculating the mean embedding from multiple word vectors is still a solid approach to the inner workings of current sentence models.
@@
Averaging is used in various scenarios. Some authors employ it to derive static embeddings from contextualized models. In [49], word vectors are learned by predicting them from the average of all contextual embeddings of words (except the target word) returned by the BERT encoder. Other authors of [53] derived static word embeddings from contextualized transformer representations by averaging word embeddings for each word in 100k different contexts. Such derived static versions are of better quality than classic Word2Vec and share their benefits, such as having tens of millions of times lower computational cost than using standard contextual embedding models [49]. The other scenario is to compose a new more expressive meta-embedding from different models and domains. As a means of composition, the authors of [54] proposed concatenation with some trainable methods. However, a later work [55] found that averaging word vectors and padding them with zeros to compensate for dimensionality mismatches is a surprisingly effective method.
@@
Despite its popularity, averaging has inherent deficiencies. The most prominent is the loss of word order information. For this reason, it is sometimes called a bag-of-words representation (that is, imagine that words lay in the bag in no specific order). As shown in [44], models based on averaging are weaker in tasks that require syntax information. The mean pooled static representations of the words “dog chased human” and “human chased dog” will be the same. In addition, simple summation can cause destructive interference, affecting valuable information. This is especially relevant in long documents. In addition to neglected sentence structure and word order, the authors of [43] also note that individual word identities are lost and noninformative words are more prominently represented than essential ones.